OpenStudy (magbak):

I WILL AWARD MEDAL!!!!!!!!!!!!!!!!!!! PLEASE HELP!!!!!!!!!!!!!!!!! A logarithmic function is the inverse of an exponential function. always sometimes never

OpenStudy (magbak):

@whpalmer4

OpenStudy (anonymous):

always

OpenStudy (uri):

Always

OpenStudy (magbak):

Ok thank you both.

OpenStudy (whpalmer4):

Yes

OpenStudy (magbak):

Thank you @whpalmer4 @uri @denis1206

OpenStudy (anonymous):

Always, because they cancel when used together. Maybe this site can help explain it better

OpenStudy (whpalmer4):

\[b^{\log_b x} = x\]

OpenStudy (magbak):

Thank you so much. I will put that formula in mind because I like having formulas because formulas always work.

OpenStudy (whpalmer4):

So, for example, \[\log_{10} 100 = 2, \text{ because } 10^2 = 100\]\[10^{\log_{10} 100} = 10^2 = 100\]The exponential function which is the inverse to the common logarithm (base 10) is \(10^x\)